OpenStudy (anonymous):

\sqrt{\frac{15 x}{x^2-196}} What is the domain?

OpenStudy (mathmale):

Do you mean this? \[\sqrt{\frac{ 15x }{ x^2-196 }}?\]

OpenStudy (mathmale):

If so, rewrite the above, EXCEPT please factor the denominator.

OpenStudy (anonymous):

Yes i mean that!

OpenStudy (mathmale):

Remember that the quantity under the radical sign can NOT be negative, although it can be either zero or positive. Note that the denominator MUST NOT = 0. Does this help towards your identification of the domain? What does domain mean, by the way?

OpenStudy (anonymous):

x value

OpenStudy (anonymous):

So x= 14?

OpenStudy (mathmale):

"domain" involves more than just "x-value." When you are asked for the "domain," you are being asked which x-values are acceptable and which are not. For example, the square root of -1 is imaginary, which would not be acceptable if you were to find only real roots.

OpenStudy (anonymous):

On my calc. it say -13 through 0

OpenStudy (mathmale):

\[\frac{ 15x }{ x^2-196 }\]

OpenStudy (mathmale):

has to be zero or greater. Notice how I've dropped the square root operator? With that in mind, what are you going to do next?

OpenStudy (mathmale):

What were y ou looking at / what were you looking for ... when you say, "on my calculator it says -3 through 0" ?

OpenStudy (anonymous):

I graphed it and then went to the table.

OpenStudy (mathmale):

Please be clear about what you're looking for when you do something like that. Once again: Your \[\frac{ 15x }{ x^2-196 }\]

OpenStudy (mathmale):

must be equal to or greater than zero (0). Why? what would you do to solve that? Your "-13 through 0" is on the right track, but what does "-13 through 0" mean? Want you to understand what you are doing and why.

OpenStudy (mathmale):

Solve:\[\frac{ 15x }{ x^2-196 }\ge0\]

OpenStudy (mathmale):

Hints: 1) First look at that numerator: 15x. At what value of x does the numerator change signs? 2) Factor the denominator. Then set each of the factors = to 0 separately. Solve each equation for x. what do you get?

OpenStudy (mathmale):

You should be coming up with three x-values here. What are they?

OpenStudy (mathmale):

OpenStudy says you are "offline." Did you lose your Internet connection, or did you simply leave without letting me know?